In a class, there are three sections, each including 20 students. In first section there are 10 boys and 10 girls, in second section, there are 15 boys and 5 girls and in the third section, there are 12 boys and 8 girls. Five students are selected from each group to form a committee of 15 students. What is the probability that all 15 students selected are girls?
1
Expert's answer
2020-04-20T13:33:43-0400


Total number of methods to select 5 students= "\binom{20}{5}=\frac{20!}{5!15!}=15504"


In the first section there are 10 boys and 10 girls,

Number of methods to select 5 girls and 0 boys= "\binom{10}{5}*\binom{10}{0}=\frac{10!}{5!5!}*1=252" ,

probability of selecting 5 girls and 0 boys "(P_1)=\frac{252}{15504}"


In the second section there are 15 boys and 5 girls,

Number of methods to select 5 girls and 0 boys="\binom{5}{5}*\binom{10}{0}=1*1=1"

probability of selecting 5 girls and 0 boys "(P_2)=\frac{1}{15504}"


In the third section there are 12 boys and 8 girls,

Number of methods to select 5 girls and 0 boys="\binom{8}{5}*\binom{12}{0}=\frac{8!}{5!3!}*1=56"

probability of selecting 5 girls and 0 boys"(P_3)=\frac{56}{15504}"


probability of all 15 students being girls="P_1*P_2*P_3"

probability of all 15 students being girls="\frac{252}{15504}*\frac{1}{15504}*\frac{56}{15504}"

probability of all15 students being girls="\bold{3.7866*10^{-9}}"